# Optimal Coupled Sensor Placement and Path-Planning in Unknown Time -Varying Environments

### v1

@@ 3.3 CRMI Optimization
See the Appendix. ∎
1:  Set $\boldsymbol{q}_{\mathrm{gr}}=\emptyset$.
@@
3:     Set $\boldsymbol{q}:=\{N_{\mathrm{G}}\}\backslash\boldsymbol{q}_{\mathrm{gr}}.$
4:     Calculate $q^{*}:=\arg\max_{q\in\boldsymbol{q}}I(J;\boldsymbol{z}(\boldsymbol{q})).$
5:     $\boldsymbol{q}_{\mathrm{gr}}=\boldsymbol{q}_{\mathrm{gr}}\cup q^{*}.$
Figure 2: Greedy CRMI optimization.
@@ 3.6 Convergence of the Proposed CSCP Algorithm
If $A$ is Hurwitz, then the CSCP algorithm terminates in a finite number of iterations.
@@
If all modes of $A$ are stable, then the pair $(A,C_{k}(\boldsymbol{q}))$ is uniformly detectable for any $\boldsymbol{q}.$ Furthermore, because $Q=\sigma_{P}^{2}\boldsymbol{I}_{N_{\mathrm{P}}},$ the pair $(A,Q)$ is controllable. By [43] it follows that the estimation error is exponentially stable. Consequently, for any $\varepsilon>0$ there is a finite iteration number $M\in{\rm Nature}$ such that $\mathrm{tr}(P_{k})\leqslant\varepsilon,$ after which the CSCP algorithm terminates. ∎
@@
Whereas Proposition 3 is sufficient, it is not necessary for the convergence of the CSCP algorithm. A less restrictive criterion for convergence, which does not assume $A$ to be Hurwitz, is discussed below.
@@
Note that this criterion cannot be verified a priori as it depends on the estimation error convariances computed during the algorithm’s execution.
@@ 4 Results and Discussion
In this section, we first provide an illustrative example of the proposed CSCP-CRMI method. Second, we compare the proposed method against a decoupled method that finds optimal sensor placement using the standard (path-independent) MI; for brevity we call this decoupled method CSCP-SMI. Third, we study the effects of varying numbers of sensors, threat parameters, and grid points on the CSCP-CRMI method. Fourth, we perform comparative study between the CSCP-CRMI and greedy sensor placements, and observe the equivalency as well as differences between the two approaches. Finally, we conduct a comparative study between the CSCP method with and without the sensor reconfiguration cost. All simulations are performed within a square workspace $\mathcal{W}=[-1,1]\times[-1,1]$ using non-dimensional units.
@@ 4.1 Illustrative Example
The implementation of CSCP-CRMI algorithm on an illustrative example is shown in Fig. 4. The number of threat parameters, grid points, and sensors are $N_{\mathrm{P}}=25$, $N_{\mathrm{G}}=49$, and $N_{\mathrm{S}}=2$, respectively. The threat parameters $N_{\mathrm{P}}$, indicated by the black dots and numbered from 1 to 25, are uniformly spaced in the workspace. The white dots represent the grid points. The initial and the goal points are represented by the bottom left and the top right grid points in the map. The evolution of the threat field estimate $\widehat{c}$ for different time steps, namely $k=1,5,11$, and $15$ is shown by a colormap. The path $\boldsymbol{v}^{*}_{k}$ of minimum estimated cost is indicated by red circles, and the sensor placement $\boldsymbol{q}_{k}$ is shown by white circles, as illustrated in the Fig. 4(a)-(d). For a specified threshold $\varepsilon=0.1$, the algorithm terminates at $k=15$ iterations, and the optimal path $\boldsymbol{v}^{*}$ is achieved.
@@ 4.2 Comparison of CSCP-CRMI and CSCP-SMI
The white regions in Fig. 6 (a) with high error covariance represents the areas where few if any sensors are placed throughout the execution of CSCP-CRMI. The darker regions represent areas where sensors were placed to reduce the estimation error covariance values orders of magnitude below the white-colored regions. Compare Fig. 6 to the optimal path found in Fig. 4(d), and we find that the CSCP-CRMI sensor placement is such that areas close to the optimal path are those with low estimation error covariance. Note, crucially, the novelty of this approach: the optimal path is at first unknown, and that the sensor placement and path-planning are performed iteratively to arrive at these results.
@@
An intensity map showing the mutual information values for each grid points from an illustrative example is shown in Fig. 7. Notably, these values are derived with consideration for only a single sensor. The brown regions represent the areas with higher CRMI values. It can be observed in Fig. 7 (a) that the CRMI regions are more visible around the vicinity of the path obtained in Fig. 4(d). Sensor is placed at the grid point, here number 49, with a maximum CRMI value. Similarly, Fig. 7(b) shows the SMI map representing the mutual information values between the state and measurement variables. The higher CRMI regions are observed around the location of the threat parameter, and for this example, sensor is placed at 7 numbered grid point.
@@
Figure 8 shows a comparison between the path cost variance $\textrm{Var}[{J}(\boldsymbol{v}^{*}_{k})]$ of the two methods. Note that for $\varepsilon=0.1$, the CSCP-SMI algorithm requires 39 iterations for convergence, which is 160% larger than the number of iterations required for CSCP-CRMI. Figure 9 shows similarly large differences in convergence rate for different numbers of sensors.
Figure 10(a) shows the variation of path cost variance with varying number of sensors. For a specified number of threat parameters $N_{\mathrm{P}}=49$ and the grid points $N_{\mathrm{G}}=49$, better convergence is achieved with more number of sensors. Using a single sensor will require a relatively large number of iterations for convergence.
@@ 4.3 Analysis of Greedy Optimization
A comparative example of sensor placements for $N_{\mathrm{P}}=25,N_{\mathrm{G}}=25$, and $N_{\mathrm{S}}=4$ using greedy and non-greedy (true optimal) is shown in Fig. 11(a). Red circles in the figure indicate sensor positions with greedy placement, whereas the black circles indicate true optimal. Note that the two sensor locations in the grid space are common for both criteria. This is a specific example where greedy optimization results in the true optimal sensor configuration. In general, greedy optimization results in near-optimal configurations.
@@
Figure 11: Comparison between greedy and non greedy CSCP-CRMI process.
@@ 4.5 Computational Complexity
The computational complexity of the Algorithm 1 depends on the complexity of CRMI optimization in Line 5, and the number of calls performed in Dijkstra’s algorithm in Line 9. CRMI calculation and maximization involves determinant and inverse computations, thus has a time complexity of $\mathcal{O}(N_{\mathrm{S}}^{3})$. By comparison, the CRMI calculation for the greedy optimization method has a complexity of $\mathcal{O}(N_{\mathrm{S}})$. The worst-case time complexity of Dijkstra’s algorithm implemented with a Fibonacci heap is $\mathcal{O}(N_{\mathrm{G}}+N_{\mathrm{G}}\log(N_{\mathrm{G}}))$ [34]. The complexity of CRMI optimization depends on the specific algorithm chosen.
@@ 5 Conclusions
In this paper, we discussed a new measure of information gain for optimal sensor placement to capture coupling of the sensor placement problem with a path-planning problem. This measure, which we call the context-relevant mutual information (CRMI), addresses the reduction in uncertainty in the path cost, rather than the environment state. We presented a coupled sensor placement and path-planning algorithm that iteratively places sensors based on CRMI maximization, updates the environment threat estimate, and then plans paths with minimum expected cost. Crucially, we showed CRMI to be a submodular function, due to which we can apply greedy optimization to arrive at near-optimal results while maintaining computational efficiency. We performed a comparative study between CSCP-CRMI and a decoupled CSCP-SMI method. The CSCP-SMI method places sensors by maximizing the standard (path-independent) mutual information of the measurments and threat state. We showed via numerical simulation examples that the CSCP-CRMI algorithm converges in less than half as many iterations compared to CSCP-SMI algorithm, which indicates a significant reduction in the number of sensor observations needed to find near-optimal paths. We also introduced and analyzed a modified cost function that addresses costs associated with distances traveled by sensors (i.e., sensor reconfiguration cost) after each iteration of placement.
@@ Acknowledgments
Consider two subsets of the sensor configurations, $\mathcal{A}$ and $\mathcal{B}$, such that $\mathcal{A}\subseteq\mathcal{B}\subseteq\boldsymbol{q}$, where $\boldsymbol{q}\subset\{N_{\mathrm{G}}\}.$ By (13), we write $I({J};\mathcal{A})=H(\mathcal{A})-H(\mathcal{A}\mid J)$ and $I({J};\mathcal{B})=H(\mathcal{B})-H(\mathcal{B}\mid J)$. For any $\𝑥\𝒒x\in\boldsymbol{q}\backslash\mathcal{B}$, the marginal CRMI gain due to a sensor placed at $x$ to in addition to those in sets $\mathcal{A}$ and $\mathcal{B}$ is
@@
where $\mathcal{S}\subseteq\{N_{\mathrm{S}}\}$ is a subset of sensor indices. Consider two subsets $\mathcal{A}$ and $\mathcal{B}$, such that $\mathcal{A}\subseteq\mathcal{B}\subseteq\{N_{\mathrm{S}}\}$ and $x\in\{N_{\mathrm{S}}\}\backslash\mathcal{B}$. In the rest of this proof, the symbol $j$ is an index over $\{N_{\mathrm{S}}\},$ i.e., $j\in\{N_{\mathrm{S}}\},$ which we avoid writing explicitly for notational convenience. Without loss of generality, we assume $\alpha_{2}=1.$
@@
An augmented state variable $\boldsymbol{\Theta}_{k-1}^{a}\in[N_{\boldsymbol{\Theta}}+N_{\boldsymbol{\omega}}+N_{\boldsymbol{\eta}}]$ is defined as $\boldsymbol{\Theta}_{k-1}^{a}=[\boldsymbol{\Theta}^{T}_{k-1}\quad\boldsymbol{\omega}^{T}_{k-1}\quad\boldsymbol{\eta}^{T}_{k-1}]^{T}$ and $12N+$ 1 sigma vectors are initialized using augmented estimated states $\widehat{\boldsymbol{\Theta}}_{k-1}^{a}$ and error covariance $P^{a}_{k-1}$ as:

### v2

@@ 3.3 CRMI Optimization
Refer to Appendix A. ∎
1: Set $\boldsymbol{q}_{\mathrm{gr}}=\emptyset$;
@@
3:  Set $\boldsymbol{q}:=\{N_{\mathrm{G}}\}\backslash\boldsymbol{q}_{\mathrm{gr}}$;
4:  Calculate $q_{i}^{*}:=\arg\max_{q\in\boldsymbol{q}}I(J;\boldsymbol{z}(\boldsymbol{q}))$;
5:  $\boldsymbol{q}_{\mathrm{gr}}=\boldsymbol{q}_{\mathrm{gr}}\cup q_{i}^{*}$.
Algorithm 2 Greedy CRMI Optimization
@@ 3.6 Convergence of the Proposed CSCP Algorithm
If $A$ is Schur, then the CSCP algorithm terminates in a finite number of iterations.
@@
If all modes of $A$ are stable, then the pair $(A,C_{k}(\boldsymbol{q}))$ is uniformly detectable for any $\boldsymbol{q}.$ Furthermore, because $Q=\sigma_{P}^{2}\boldsymbol{I}_{N_{\mathrm{P}}},$ the pair $(A,Q)$ is controllable. By [50], it follows that the estimation error is exponentially stable. Consequently, for any $\varepsilon>0$ there is a finite iteration number $M\in{\rm Nature}$ such that $\mathrm{tr}(P_{k})\leqslant\varepsilon,$ after which the CSCP algorithm terminates. ∎
@@
Whereas Proposition 3 is sufficient, it is not necessary for the convergence of the CSCP algorithm. A less restrictive criterion for convergence, which does not assume $A$ to be Schur, is discussed below.
@@
Note that this criterion cannot be verified a priori as it depends on the estimation error covariances computed during the algorithm’s execution.
@@ 4 Results and Discussion
In this section, we first provide an illustrative example of the proposed CSCP-CRMI method. Second, we compare the proposed method against a decoupled method that finds optimal sensor placement using the standard (path-independent) MI; for brevity we call this decoupled method CSCP-SMI. Third, we study the effects of varying numbers of sensors, threat parameters, and grid points on the CSCP-CRMI method. Fourth, we perform comparative study between the CSCP-CRMI and greedy sensor placements, and observe the equivalency as well as differences between the two approaches. Finally, we conduct a comparative study between the CSCP method with and without the sensor reconfiguration cost. All simulations are performed within a square workspace $\mathcal{W}=[-1,1]\times[-1,1]$ using non-dimensional units in a Cartesian coordinate axes system.
A MATLAB®-based implementation of the CSCP method used for producing these results is available at this repository: https://github.com/prakashpoudel2014/CSCP_time_varying.
@@ 4.1 Illustrative Example
(a) Error ($k=1$)
(b) Error ($k=5$)
(c) Error ($k=11$)
(d) Error ($k=15$)
Figure 4: Error between the mean estimate and ground truth during CSCP-CRMI, for $N_{\mathrm{P}}=25$ and $N_{\mathrm{G}}=49$.
{subfigmatrix}
2
@@
The implementation of CSCP-CRMI algorithm on an illustrative example is shown in Fig. 3. The number of threat parameters, grid points, and sensors are $N_{\mathrm{P}}=25$, $N_{\mathrm{G}}=49$, and $N_{\mathrm{S}}=2$, respectively. The threat parameters $N_{\mathrm{P}}$, indicated by the black dots and numbered from 1 to 25, are uniformly spaced in the workspace. The white dots represent the grid points. The initial and the goal points are represented by the bottom left and the top right grid points in the map. The evolution of the threat field estimate $\widehat{c}$ for different time steps, namely $k=1,5,11$, and $15$ is shown by a colormap. The path $\boldsymbol{v}^{*}_{k}$ of minimum estimated cost is indicated by red circles, and the sensor placement $\boldsymbol{q}_{k}$ is shown by white circles, as illustrated in the Fig. 3(a)-(d). For a specified threshold $\varepsilon=0.1$, the algorithm terminates at $k=15$ iterations, and the optimal path $\boldsymbol{v}^{*}$ is achieved. Fig. 4(a)-(d) show the absolute error between the true threat field $c$ and the estimated field $\widehat{c}$, computed as |cc^||c-$\widehat${c}|, for the respective iterations $k=1,5,11,15$. These plots illustrate how the estimation accuracy improves as the algorithm progresses.
@@ 4.2 Comparison of CSCP-CRMI and CSCP-SMI
The white regions in Fig. 6(a) with high error covariance represent areas where few, if any, sensors are placed throughout the execution of CSCP-CRMI. The darker regions represent areas where sensors were placed to reduce the estimation error covariance values orders of magnitude below those of the white-colored regions. Compare Fig. 6 to the optimal path found in Fig. 3(d), and we find that the CSCP-CRMI sensor placement is such that areas around the optimal path generally have lower estimation error covariance. Note that the path does not exclusively follow regions of minimal covariance. The planning objective is to minimize cumulative threat exposure, the optimal trajectory occasionally favors areas with slightly higher estimation uncertainty if the expected threat intensity in those areas is substantially lower. Note, crucially, the novelty of this approach: the optimal path is at first unknown, and that the sensor placement and path-planning are performed iteratively to arrive at these results.
@@
An intensity map showing the mutual information values for each grid points from an illustrative example is shown in Fig. 7. Notably, these values are derived with consideration for only a single sensor. The brown regions represent the areas with higher CRMI values. It can be observed in Fig. 7 (a) that the CRMI regions are more visible around the vicinity of the path obtained in Fig. 3(d). The optimal path, indicated by blue circles, is overlaid on the CRMI intensity map for clarity. Sensor is placed at the grid point, here number 49, with a maximum CRMI value. Similarly, Fig. 7(b) shows the SMI map representing the mutual information values between the state and measurement variables. The higher CRMI regions are observed around the location of the threat parameter, and for this example, sensor is placed at 7 numbered grid point.
@@
Figure 8 shows a comparison between the path cost variance $\textrm{Var}[{J}(\boldsymbol{v}^{*}_{k})]$ of the two methods. Note that for $\varepsilon=0.1$, the CSCP-SMI algorithm requires 39 iterations to converge, whereas the CSCP-CRMI algorithm requires only 15 iterations. This indicates that the number of iterations for the SMI based method is 160% larger than CSCP. Figure 9 shows similarly large differences in convergence rate for different numbers of sensors.
Figure 10(a) shows the variation of path cost variance with varying number of sensors. For a specified number of threat parameters, $N_{\mathrm{P}}=49$ and the grid points $N_{\mathrm{G}}=49$, better convergence is achieved with more number of sensors. Using a single sensor will require a relatively large number of iterations for convergence.
@@ 4.3 Analysis of Greedy Optimization
A comparative example of sensor placements for $N_{\mathrm{P}}=25,N_{\mathrm{G}}=25$, and $N_{\mathrm{S}}=4$ using greedy and non-greedy (true optimal) is shown in Fig. 11(a). Red circles in the figure indicate sensor positions with greedy placement, whereas the black circles indicate true optimal. Note that the two sensor locations in the grid space are common for both criteria. This is a specific example where greedy optimization results in the true optimal sensor configuration. For a submodular function, greedy optimization results in near-optimal configurations.
@@
Figure 11: Comparison between greedy and non greedy CSCP-CRMI.
@@ 4.4 Analysis of Sensor Reconfiguration Cost
Finally, although the basis functions used in the previous results had uniformly spaced centers and non-significant intersections in regions of support, the proposed CSCP method is not restricted by these assumptions. For instance, Figure 14 shows different CSCP iterations in a threat field constructed using overlapping basis functions and non-uniformly spaced centers.
(a) $k=4$
(b) $k=10$
(c) $k=15$
(d) $k=19$
Figure 14: CSCP iterations for non uniformly spaced basis functions, with $N_{P}=49$, $N_{G}=121,$ and $N_{S}=2$.
@@ 4.5 Computational Complexity
The computational complexity of the Algorithm 1 depends on the complexity of CRMI optimization in Line 5, and the number of calls performed in Dijkstra’s algorithm in Line 9. CRMI calculation and maximization involves determinant and inverse computations, thus has a time complexity of $\mathcal{O}(N_{\mathrm{S}}^{3})$. By comparison, the CRMI calculation for the greedy optimization method has a complexity of $\mathcal{O}(N_{\mathrm{S}})$. The worst-case time complexity of Dijkstra’s algorithm implemented with a Fibonacci heap is $\mathcal{O}(N_{\mathrm{G}}+N_{\mathrm{G}}\log(N_{\mathrm{G}}))$ [41]. The complexity of CRMI optimization depends on the specific algorithm chosen. CSCP addresses scenarios where sensor resources are limited. Scaling up should then be considered in the context of the grid size, which may be achieved by relaxation to a continuous workspace and a suitable optimization method, e.g., evolutionary global optimization methods, as previously mentioned. Further computational efficiency may be achieved via decentralized estimation, as we had reported earlier in [51], albeit with a naïve sensor placement method.
@@ 5 Conclusions
In this paper, we discussed a new measure of information gain for optimal sensor placement to capture coupling of the sensor placement problem with a path-planning problem. This measure, which we call the context-relevant mutual information (CRMI), addresses the reduction in uncertainty in the path cost, rather than the environment state. We presented a coupled sensor placement and path-planning algorithm that iteratively places sensors based on CRMI maximization, updates the environment threat estimate, and then plans paths with minimum expected cost. Crucially, we showed CRMI to be a submodular function, due to which we can apply greedy optimization to arrive at near-optimal results while maintaining computational efficiency. We performed a comparative study between CSCP-CRMI and a decoupled CSCP-SMI method. The CSCP-SMI method places sensors by maximizing the standard (path-independent) mutual information of the measurments and threat state. We showed via numerical simulation examples that the CSCP-CRMI algorithm converges in less than half as many iterations compared to CSCP-SMI algorithm, which indicates a significant reduction in the number of sensor observations needed to find near-optimal paths. We also introduced and analyzed a modified cost function that addresses costs associated with distances traveled by sensors (i.e., sensor reconfiguration cost) after each iteration of placement. For future work, we will consider other real-world sensor constraints such as communication limitations, energy consumption, and sensor reliability issues.
@@ Acknowledgments
[…]
@@
[…]
@@
An augmented state variable $\boldsymbol{\Theta}_{k-1}^{a}\in[N_{\boldsymbol{\Theta}}+N_{\boldsymbol{\omega}}+N_{\boldsymbol{\eta}}]$ is defined as $\boldsymbol{\Theta}_{k-1}^{a}=[\boldsymbol{\Theta}^{T}_{k-1}\quad\boldsymbol{\omega}^{T}_{k-1}\quad\boldsymbol{\eta}^{T}_{k-1}]^{T}$ and $2N+1$ sigma vectors are initialized using augmented estimated states $\widehat{\boldsymbol{\Theta}}_{k-1}^{a}$ and error covariance $P^{a}_{k-1}$ as: